OpenStudy (anonymous):

A parallel-plate capacitor is connected to a battery that maintains a constant potential difference between the plates. If the spacing between the plates is doubled, how is the magnitude of charge on the plates affected?

OpenStudy (sidsiddhartha):

for parallel plate capacitors capacitence is \[C=\frac{ e_0*A }{ d }\] where e_0=permittivity of the medium between the plates A=area d=distence between the plates

OpenStudy (sidsiddhartha):

again we know that simply \[C=Q/V\] ok??

OpenStudy (anonymous):

they are asking to give the answer as Qnew/Qold

OpenStudy (anonymous):

that is what is confusing me. usually they ask for an answer "in terms of" but i just have a problem figuring out how to answer the question

OpenStudy (sidsiddhartha):

so we've got two equations \[C=\frac{ e_0*A }{ d}........................(1)\]and \[C=\frac{ Q }{ V }.......................................(2)\]now equate them \[\frac{ E_0*A }{ d }=\frac{ Q }{V}\] so from this equation we can see that Q=charge is inversely proportional to the distence between the plates\[Q \alpha (1/d)\] \[\frac{ Q_(new) }{ Q_(old) }=\frac{ d_{old} }{ d_{new} }\]

OpenStudy (sidsiddhartha):

got this?

OpenStudy (anonymous):

let me test it out!

OpenStudy (anonymous):

it said that was not correct

OpenStudy (anonymous):

Incorrect; Try Again; 3 attempts remaining The correct answer does not depend on: dnew, dold.

OpenStudy (sidsiddhartha):

no it will surely not depend upon dnew and dold u have to substitute values of them \[\frac{ Q_{new} }{ Q_{old}}=\frac{ d_{old} }{ 2*d_{old}}=1/2\]

OpenStudy (sidsiddhartha):

so \[Q_{new}=\frac{ Q_{old} }{ 2}\]

OpenStudy (sidsiddhartha):

now try this

OpenStudy (anonymous):

i entered .5 and it was correct

OpenStudy (anonymous):

tricky!!! I hate it when I cant figure out what terms to answer a question in!

OpenStudy (sidsiddhartha):

lol :)